# A Geometric Tension Dynamics Model of Epithelial Convergent Extension

### v2

@@ 2 Additional simulation results
In this section we report additional results from the tissue scale simulations and/or explain implementation details.
@@ 2.1 Influence of isogonal stretching on the LTC distribution
Here, we consider the influence of the reference cell shape $S_{0}$ which determines the isogonal potential via the elastic energy. As discussed in the main text, isogonally stretching or compressing cells along the axis of tension anisotropy can delay or accelerate T1s by moving the T1-threshold in LTC space. As reported in the main text, if we implement isogonal stretching by choosing an anisotropic $S_{0}$, i.e. $S_{0}=3\ell_{0}\;\mathrm{diag}(1-t,1+t),\;t\in[0,1]$, we find that that a higher bridge bias emerges at the early phase of convergent extension (Fig. 5). We chose an anisotropic reference shape $t=1/3$ to model the isogonal stretching caused by the ventral furrow before onset of GBE. In the experimental data, this isogonal stretch decays; analogously, we linearly ramp the reference shape anisotropy down so that $t=0$ at 50 minutes simulation time. We note that our modelling of isogonal stretching is limited, since it is encoded in a model parameter ($S_{0}$) instead of being created dynamically, for example by external forces applied to the boundary. Time traces of the LTC parameter in the absence of isogonal stretching are show in Fig. S8) below.
@@ 2.2 Alternative forms of positive tension feedback
The overall dynamics of this model is very similar to Eq. (5) considered in the main text. However, the tension feedback is “more aggressive” since now the total tension (triangle perimeter) increases as the tension triangle becomes more anisotropic. This mirrors a situation where in addition to the anisotropy also the total myosin levels increase during GBE [71]. Therefore, one observes slightly larger amounts of convergent extension for identical initial conditions.
@@
To integrate Eq. (S3), we use the explicit Euler-Maruyama scheme. We find that the convergent extension phenomenology is robust to low to moderate levels of noise (i.e. $\sigma<T_{ij}^{n}$). However, higher $\sigma$ leads to a final LTC distribution with somewhat more anisotropic triangles, indicating larger disorder, and reduces the amount of total convergent extension.
@@ 2.3 Saturating tension feedback
where $T_{-}<T_{c}<T_{+}$ are the low, unstable, and high tension fixed point, respectively. We set $T_{-}=0,\;T_{c}=1$ matching the fixed points of the main feedback model Eq. (5) we consider in this manuscript. The simulations with saturating feedback correspond to $T_{+}=1.2$, the control simulations are $T_{+}=3.33$. Note that we adjusted the time step and all other rate parameters of the simulation so that $\dot{T}$ is similar across different choices of $T_{+}$.
Figure S5: Tension and length dynamics of of intercalating interface for T1 events in winner-takes-all and saturating feedback simulations. Shaded area indicates standard deviations. Time $t=0\mathrm{min}$ indicates the moment of T1 transition. Winner-takes-all feedback matches the corresponding experimental data well (see the companion paper [17]), while saturating feedback does not, featuring a rapid instead of gradual collapse and reemergence of interfaces. Interface length is measured relative to the average interface length $\overline{\ell}$. T1s occur at a non-zero length due to a numerical cutoff.
@@
We note that in the case of saturating tension feedback, a relevant fraction of T1s (approx. 20%) are reversible, i.e. the newly formed junction re-collapses, which occurs in less than 1% of cases for non-saturating feedback. Only irreversible T1s are counted in Fig. 6E. Because saturating feedback only very rarely produces reversible T1s, all quantifications other than Fig. 6 show the total T1 count without any filtering.
Reversible T1s are defined as follows. Consider an edge $ij$ between cells $i$ and $j$, the central edge in a quartet of cells $ikjl$ (in clockwise order). The collapse of edge $ij$ creates an edge $kl$. This T1 is considered reversible if $kl$ collapses in turn to give rise to a connection $ij$ again, i.e. $kl$ re-collapses before any of the outer edges $il$, $lj$, $ik$, or $ki$, of the quartet have collapsed. In this case, the local quartet topology returns to its original state. We also filter out nested sequences of multiple reversible T1s, but such events are extremely rare.
Figure S6: Final tissue shapes (40 min simulation time) for the alternative tissue dynamics models discussed in the SI. Simulations were initialized with an “ordered” tension triangulation and initial anisotropy $s=0.2$. Cells are colored by coordination number (same color code as in Fig. 2A).
@@ 2.4 Effect of hexatic order orientation
When carrying out simulations in which the tissue is initially ordered, the initial tension triangulation corresponds to a (rectangular) patch of a triangular lattice. We can choose two possible orientations of the lattice plane with respect to the tension anisotropy – one creating initially mainly tension cables, one mainly bridges. This corresponds to the orientation of the hexatic order $\Psi_{6}$ w.r.t. the axis of tension anisotropy, and we refer to these orientations as “bridge” and “cable” initial conditions. We find that the qualitative behavior of the model is independent of this choice, and positive tension feedback is capable of creating coherently ordered T1s even if future collapsing edges are not singled out in the initial condition. To underline this point, Fig. 2 uses “cable” initial conditions. For the determination of LTC distribution dynamics in main text Fig. 5, as well as the heat map Fig. 3A, we pool simulations from cable and bridge initial conditions. When showing snapshots from a simulation, the “parallel” initial condition most easily creates the recognizable pattern of alternating high and low tensions reported in the companion paper.
A quantitative comparison is shown in Fig. S7. We see that an initial bridge bias leads to higher bridge fraction later on, and somewhat larger amounts of convergent extension.
Figure S8: Dependence of model behavior on parameters. Each row of plots shows the dependence on one parameter or modelling choice. The three columns follow the same convention as Fig. S7B–D. The passive relaxation time $\tau_{p}$, the balancing time $\tau_{\mathrm{balance}}$, the feedback cutoff, and the initial condition vertex noise are explained in more detail in SI Sec. 4.
@@ 2.5 Interface extension is independent of tissue extension
Our simulations support the notion that interface extension does not depend on tissue extension, even though our simulations contain no additional active mechanisms to elongate new interfaces. In the main text, we showed simulations of a tissue geometry without a soft, passive region to compensate the convergence of the active tissue (analogous to a Toll[RM9] mutant). Because of incompressibility, the active tissue also does not extend. Nevertheless, new interfaces elongate at rates very similar to control simulations (Fig. 8). We now support this conclusion with an additional simulation. Mimicking experiments where tissue extension is blocked by a cauterization fence [8], we block tissue extension by adding slip walls along the AP boundaries of the tissue patch (Fig. S9A). For numerical reasons, we add a small passive zone in front of these additional boundaries, and enforce the slip-wall condition with a softer penalty than for the dorsal slip walls. This is done to avoid direct conflicts between the boundary conditions and the angle constraints from the triangulation. In these simulations, the active tissue does not extend, and cells deform to compensate T1s (Fig. S9A). In a control simulations (Fig. 8), by 15 minutes, the tissue has elongated by 30% by the time shown. However, new interfaces still successfully extend (Fig. S9B), like in the Toll[RM9] simulations shown in Fig. 8. The overall speed of intercalations is slightly reduced because of the additional slip walls. Note that occasionally, we observe rupture of the simulated tissue from the simulation boundary because of the internal forces generated by T1 transitions. Similar rupture events are seen in experiment.
Figure S9: Interfaces extend even if tissue extension is blocked. A Snapshots from a simulation where tissue extension is blocked by slip walls along the AP boundaries of the tissue patch. This mimics experiments where the germ band is blocked from extending via cauterization of the ventral tissue close to the posterior pole, or mutations in the posterior midgut [8]. B Even though net tissue deformation is blocked, the interface length during active T1s is qualitatively similar to simulations of freely deforming tissues (cf. “WT” in Fig. 8B).
Figure S10: Elastic energy in germ-band simulations tissue-scale simulations (Fig. 7). Note that in “WT” simulations, the elastic energy does not increase over time. If no passive tissue exists to compensate the extension of the germ band, cell elongate and elastic energy builds up.
@@ 2.6 Scale separation between cortical tensions and cell shape elastic energy
As discussed in the main text, we have formulated a model of epithelial tissue mechanics in the limit where the cortical tensions are the dominant source of stress. This separation of scales is encoded by the small parameter $\varepsilon$ in Eq. (1). To lowest order in $\varepsilon$, force balance of the cortical tensions impose geometric constraints on angles under which the subdominant “cell-shape energy” is relaxed. In other words, the tension-driven and isogonal degrees of freedom are clearly separated. To higher order in $\varepsilon$, the cell-shape energy changes the angles, thus mixing the tension-driven and isogonal modes. How small does $\varepsilon$ have to be such that we can ignore this higher order effect?
To answer this question, we need to compare the tension force balance to the force on vertex angles exerted by the cell-shape energy $\varepsilon\sum_{i}E_{\mathcal{C}}(\{\mathbf{r}_{ijk}\})$ (defined in Eq. (3)). This force is given by the elastic energy gradient $\nabla E_{\mathcal{C}}=\partial_{\mathbf{r}_{ijk}}\sum_{i}E_{\mathcal{C}}(\{\mathbf{r}_{ijk}\})$ at the vertex positions $\mathbf{r}_{ijk}^{*}=\mathrm{argmin}_{\mathbf{r}_{ij}\perp\mathbf{t}_{ij}}\sum_{i}E_{\mathcal{C}}(\{\mathbf{r}_{ijk}\})$ which minimize the elastic energy under the angle constraints imposed by the tension triangulation (i.e. the tension triangulation edges $\mathbf{t}_{ij}$ and the cell edges $\mathbf{r}_{ij}$ are orthogonal). At this minimum, a displacement $\delta\mathbf{r}_{ijk}=-h\nabla E_{\mathcal{C}}(\mathbf{r}^{*})$ for a small $h>0$ has no isogonal component and acts purely on angles. We impose this displacement and calculate the tension- and cell-elasticity energy cost via the the non-dimensional ratio
@@
This expression on data from the simulation shown in Fig. 2A, we find $f\approx-3\mu\overline{\ell}/\overline{T}$, where $\overline{\ell}$ and $\overline{T}$ are the average interface length and tension, respectively. They set the basic units of length and force. In our simulations, we set $\overline{T}=1$ and ¯$\overline{\ell}=1/\sqrt{3}$ (this ensures that the cell sizes are compatible with the Voronoi dual of the tension triangulation [17]). $f$ is negative because the cell-shape energy can relax further by changing the angles, compared to when it is constraint to the isogonal (fixed-angle) subspace. Notably, only the shear modulus $\mu$ of the cells contributes, indicating that the effect of bulk elasticity on the angles is negligible. The numerical factor of $3$ is a consequence of the particular way we have defined the elastic energy and not fundamental. Given our choice of units, the parameter setting $\mu=1$, we must have $\varepsilon\lesssim 0.2$ for the separation of tension-driven and isogonal modes to be valid.
@@
To test this, we numerically ran simulations with different values of $\varepsilon$, $0.001$ (our default value), $0.01$, $0.1$, and $1$. As shown in Fig. S11, we find almost identical behavior of the simulations for $\varepsilon=0.001$ and $\varepsilon=0.1$. By contrast, for $\varepsilon=1$ tissue shape change is impaired since the tensions are insufficient to overcome the tendency of the cell-shape energy to drive cells to regular hexagon shapes. In other words, the cell-shape energy imposes energy barriers that prevent T1 transitions. In particular, we find that at $\varepsilon=1$, more than half of T1 transitions reverse, and are hence ineffective, while this is the case for less than $2\%$ of T1s at $\varepsilon=0.001$. Without explicit saturation of the positive tension feedback, this suppression of T1s leads to buildup of extreme tension anisotropy, eventually leading to numerical problems.
Figure S12: Dependence of convergent extension on initial order and anisotropy for vanishing shear rigidity. Heatmap of elongation (square root of aspect ratio change) as in main text Fig. 3A, for $\mu=0$ (low shear rigidity). In Fig. 3A, we used $\mu=1$, the default value in this paper. For lower shear rigidity, less convergent extension takes place, as discussed in Fig. 8G. The overall structure of the phase diagram remains similar, with weaker dependence on initial order. Note that the simulation was performed with free boundary conditions, i.e. there is no external resistance to tissue deformation. An adjacent passive tissue with non-zero shear modulus would suppress convergent extension further (cf. Fig. 8F). For this supplementary figure, we only carried out 3 replicate simulations.

### v3

@@ 2 Additional simulation results
In this section, we report additional results from the tissue scale simulations and/or explain implementation details.
2.1 Dynamics of the shape index
A common measure for the degree of order in a polygonal tiling is the cell shape index ($s=P/\sqrt{A}$, with cell perimeter $P$ and area $A$), shown in Fig. S5A. This shape index is high when cell shapes are elongated or irregular and approaches the minimal value $s\approx 3.72$ for a regular hexagon. The shape index has particular relevance in vertex models employing an area–perimeter elastic energy [Hufnagel.etal2007, Farhadifar.etal2007]. In these models, the target shape index is a control parameter that drives a solid-to-fluid transition [Bi.etal2015]: a high target shape index ($s_{0}>s_{0}^{*}\approx 3.81$) is associated with tissue fluidity since it allows for cell rearrangements, while a low cell target shape index gives rise to a solid state. In both cases, the observed shape index is controlled by the target shape index. In contrast, in our simulations with actively driven T1s, disorder, and thus a high observed shape index, is the consequence of cell rearrangements, rather than their cause. Notably, we find more tissue flow when the observed shape index is initially low (Fig. S5A). The question of the solid vs fluid character of tissue is addressed in more detail in the last part of the results section and in the discussion.
2.2 Dynamics in order-anisotropy-extension phase space
As we have seen in the main text, both order and anisotropy decrease as cells rearrange (cf. Fig. 2E). When plotting the remaining extent of convergent extension $\sqrt{a_{\mathrm{final}}/a(t)}$ against $p_{6}(t)$ and $||\langle\sqrt{2}\tilde{Q}(t)\rangle||$, we find trajectories that approximately lie in a common plane and converge to a fixed point at vanishing anisotropy $||\langle\sqrt{2}\tilde{Q}\rangle||\approx 0$ and $p_{6}\approx 0.4$ (Fig. S5B). Based on these results, we hypothesize an empirical law for feedback-driven convergent extension, based on the instantaneous hexagon fraction and tension anisotropy:
$\sqrt{\frac{a_{\mathrm{final}}}{a(t)}}\approx c_{Q}||\langle\sqrt{2}\tilde{Q}(t)\rangle||+c_{6}p_{6}(t)$ (S2)
With the coefficients determined by a linear fit, (S2) predicts the remaining extent of convergent extension with less than 10% mean absolute error (Fig. S5C). The empirical relation (S2) is specific to the particular choice of microscopic tension dynamics employed in the simulations: a systematic search for a more general relation linking microscopic ($p_{6},||\langle\sqrt{2}\tilde{Q}(t)\rangle||$) and macroscopic ($a(t)$) quantities is beyond the scope of the present work.
@@ 2.3 Influence of isogonal stretching on the LTC distribution
Here, we consider the influence of the reference cell shape $S_{0}$ which determines the isogonal potential via the elastic energy. As discussed in the main text, isogonal stretching or compression of cells along the axis of tension anisotropy can delay or accelerate T1s by moving the T1-threshold in LTC space. As reported in the main text, if we implement isogonal stretching by choosing an anisotropic $S_{0}$, i.e. $S_{0}=3\ell_{0}\;\mathrm{diag}(1-t,1+t),\;t\in[0,1]$, we find that a higher bridge bias emerges at the early phase of convergent extension (Fig. 4). We chose an anisotropic reference shape $t=1/3$ to model the isogonal stretching caused by the ventral furrow before the onset of GBE. In the experimental data, this isogonal stretch decays; analogously, we linearly ramp the reference shape anisotropy down so that $t=0$ at 50 minutes simulation time. We note that our modeling of isogonal stretching is limited since it is encoded in a model parameter ($S_{0}$) instead of being created dynamically, for example by external forces applied to the boundary. Time traces of the LTC parameter in the absence of isogonal stretching are shown in Fig. S9) below.
@@ 2.4 Alternative forms of positive tension feedback
The overall dynamics of this model are very similar to (5) considered in the main text. However, the tension feedback is “more aggressive” since now the total tension (triangle perimeter) increases as the tension triangle becomes more anisotropic. This mirrors a situation where in addition to the anisotropy also the total myosin levels increase during GBE [Lefebvre.etal2022]. Therefore, one observes slightly larger amounts of convergent extension for identical initial conditions.
@@
To integrate (S4), we use the explicit Euler-Maruyama scheme. We find that the convergent extension phenomenology is robust to low to moderate levels of noise (i.e. $\sigma<T_{ij}^{n}$). However, higher $\sigma$ leads to a final LTC distribution with somewhat more anisotropic triangles, indicating larger disorder, and reduces the amount of total convergent extension.
@@ 2.5 Saturating tension feedback
where $T_{-}<T_{c}<T_{+}$ are the low, unstable, and high tension fixed points, respectively. We set $T_{-}=0,\;T_{c}=1$ matching the fixed points of the main feedback model (5) we consider in this manuscript. The simulations with saturating feedback correspond to $T_{+}=1.2$, the control simulations are $T_{+}=3.33$. Note that we adjusted the time step and all other rate parameters of the simulation so that $\dot{T}$ is similar across different choices of $T_{+}$.
@@
For saturating tension feedback, we observe that approximately 20% of T1s reverse, i.e. the newly formed junction re-collapses. For winner-takes-all feedback, less than 1% of T1s reverse. Reversing T1s are defined as follows. Consider an edge $ij$ between cells $i$ and $j$, the central edge in a quartet of cells $ikjl$ (in clockwise order). The collapse of edge $ij$ creates an edge $kl$. This T1 is considered reversible if $kl$ collapses, in turn, to give rise to a connection $ij$ again, i.e. $kl$ re-collapses before any of the outer edges $il$, $lj$, $ik$, or $ki$, of the quartet have collapsed. In this case, the local quartet topology returns to its original state. We also filter out nested sequences of multiple reversible T1s, but such events are extremely rare.
@@ 2.6 Effect of hexatic order orientation
When carrying out simulations in which the tissue is initially ordered, the initial tension triangulation corresponds to a (rectangular) patch of a triangular lattice. We can choose two possible orientations of the lattice plane with respect to the tension anisotropy – one creating initially mainly tension cables, one mainly bridges. This corresponds to the orientation of the hexatic order $\Psi_{6}$ w.r.t. the axis of tension anisotropy, and we refer to these orientations as “bridge” and “cable” initial conditions. We find that the qualitative behavior of the model is independent of this choice, and positive tension feedback is capable of creating coherently ordered T1s even if future collapsing edges are not singled out in the initial condition. To underline this point, Fig. 2 uses “cable” initial conditions. For the determination of LTC distribution dynamics in main text Fig. 4, as well as the heat map Fig. 2D, we pool simulations from cable and bridge initial conditions. When showing snapshots from a simulation, the “parallel” initial condition most easily creates the recognizable pattern of alternating high and low tensions reported in the companion paper.
A quantitative comparison is shown in Fig. S8. We see that an initial bridge bias leads to a higher bridge fraction later on, and somewhat larger amounts of convergent extension.
@@ 2.7 Interface extension is independent of tissue extension
Our simulations support the notion that interface extension does not depend on tissue extension, even though our simulations contain no additional active mechanisms to elongate new interfaces. In the main text, we showed simulations of a tissue geometry without a soft, passive region to compensate for the convergence of the active tissue (analogous to a Toll[RM9] mutant). Because of incompressibility, the active tissue also does not extend. Nevertheless, new interfaces elongate at rates very similar to control simulations (Fig. S10B). We further support this conclusion with an additional simulation. Mimicking experiments where tissue extension is blocked by a cauterization fence [Collinet.etal2015], we block tissue extension by adding slip walls along the AP boundaries of the tissue patch (Fig. S10A). For numerical reasons, we add a small passive zone in front of these additional boundaries and enforce the slip-wall condition with a softer penalty than for the dorsal slip walls. This is done to avoid direct conflicts between the boundary conditions and the angle constraints from the triangulation. In these simulations, the active tissue does not extend, and cells deform to compensate T1s (Fig. S10A). However, new interfaces still successfully extend (Fig. S10B), like in the Toll[RM9] simulations shown in Fig. 7. The overall speed of intercalations is slightly reduced because of the additional slip walls. Note that occasionally, we observe rupture of the simulated tissue from the simulation boundary because of the internal forces generated by T1 transitions. Similar rupture events are seen in experiments.
@@ 2.8 Scale separation between cortical tensions and cell shape elastic energy
As discussed in the main text, we have formulated a model of epithelial tissue mechanics in the limit where the cortical tensions are the dominant source of stress. This separation of scales is encoded by the small parameter $\varepsilon$ in (1). To lowest order in $\varepsilon$, force balance of the cortical tensions imposes geometric constraints on angles under which the subdominant “cell-shape energy” is relaxed. In other words, the tension-driven and isogonal degrees of freedom are clearly separated. To higher order in $\varepsilon$, the cell-shape energy changes the angles, thus mixing the tension-driven and isogonal modes. How small does $\varepsilon$ have to be such that we can ignore this higher-order effect?
To answer this question, we need to compare the tension force balance to the force on vertex angles exerted by the cell-shape energy $\varepsilon\sum_{i}E_{\mathcal{C}}(\{\mathbf{r}_{ijk}\})$ (defined in (3)). This force is given by the elastic energy gradient $\nabla E_{\mathcal{C}}=\partial_{\mathbf{r}_{ijk}}\sum_{i}E_{\mathcal{C}}(\{\mathbf{r}_{ijk}\})$ at the vertex positions $\mathbf{r}_{ijk}^{*}=\mathrm{argmin}_{\mathbf{r}_{ij}\perp\mathbf{t}_{ij}}\sum_{i}E_{\mathcal{C}}(\{\mathbf{r}_{ijk}\})$ which minimize the elastic energy under the angle constraints imposed by the tension triangulation (i.e. the tension triangulation edges $\mathbf{t}_{ij}$ and the cell edges $\mathbf{r}_{ij}$ are orthogonal). At this minimum, a displacement $\delta\mathbf{r}_{ijk}=-h\nabla E_{\mathcal{C}}(\mathbf{r}^{*})$ for a small $h>0$ has no isogonal component and acts purely on angles. We impose this displacement and calculate the tension- and cell-elasticity energy cost via the non-dimensional ratio
@@
This expression on data from the simulation shown in Fig. 2A, we find $f\approx-3\mu\overline{\ell}/\overline{T}$, where $\overline{\ell}$ and $\overline{T}$ are the average interface length and tension, respectively. They set the basic units of length and force. In our simulations, we set $\overline{T}=1$ and ¯$\overline{\ell}=1/\sqrt{3}$ (this ensures that the cell sizes are compatible with the Voronoi dual of the tension triangulation [Brauns.etal2024elife]). $f$ is negative because the cell-shape energy can relax further by changing the angles, compared to when it is constrained to the isogonal (fixed-angle) subspace. Notably, only the shear modulus $\mu$ of the cells contributes, indicating that the effect of bulk elasticity on the angles is negligible. The numerical factor of $3$ is a consequence of the particular way we have defined the elastic energy and is not fundamental. Given our choice of units, the parameter setting $\mu=1$, we must have $\varepsilon\lesssim 0.2$ for the separation of tension-driven and isogonal modes to be valid.
@@
To test this, we numerically ran simulations with different values of $\varepsilon$, $0.001$ (our default value), $0.01$, $0.1$, and $1$. As shown in Fig. S11, we find almost identical behavior of the simulations for $\varepsilon=0.001$ and $\varepsilon=0.1$. By contrast, for $\varepsilon=1$ tissue shape change is impaired since the tensions are insufficient to overcome the tendency of the cell-shape energy to drive cells to regular hexagon shapes. In other words, the cell-shape energy imposes energy barriers that prevent T1 transitions. In particular, we find that at $\varepsilon=1$, more than half of T1 transitions reverse, and are hence ineffective, while this is the case for less than $2\%$ of T1s at $\varepsilon=0.001$. Without explicit saturation of the positive tension feedback, this suppression of T1s leads to a buildup of extreme tension anisotropy, eventually leading to numerical problems.
3 T1-threshold in LTC space
To find the resulting T1 threshold in the LTC space, we need to express the tension $T$ on the cell quartet’s inner interface in terms of the angles $\alpha,\alpha^{\prime},\beta,\beta^{\prime}$. Reasoning that on average the two tension triangles in a quartet will be similar, we consider the simplified case $\alpha=\alpha^{\prime},\beta=\beta^{\prime}$. Normalizing by the average tension we find $T=3\sin(\beta)/(\sin\alpha+\sin\beta+\sin(\alpha+\beta))$. Since the problem is now reduced to the shape of one triangle, the T1 condition $\ell_{\mathrm{ref}}(\alpha,\beta)+\Delta\ell_{\mathrm{iso}}=0$ defines a threshold line in the LTC space. Figure 3E shows the T1 threshold as a function of the isogonal strain $\Delta\ell_{\mathrm{iso}}/\ell_{0}$. The critical tension anisotropy drops to zero as $\Delta\ell_{\mathrm{iso}}/\ell_{0}$ approaches $-1$, the critical isogonal strain for purely passive T1s which takes place for isotropic tension (i.e. equilateral tension triangles). Vice versa, positive isogonal strain shifts the T1 threshold to a higher magnitude of tension anisotropy. In principle, the above geometric reasoning can be generalized to an arbitrary tension “kite” composed of two different tension triangles. However, the shape space of such kites is four-dimensional (since there are four independent angles $\alpha,\alpha^{\prime},\beta,\beta^{\prime}$) which precludes the intuitive visualization that the single-triangle LTC space provides.